# BuildingView: Constructing Urban Building Exteriors Databases with Street View Imagery and Multimodal Large Language Model

### v1

@@ 3.3. Annotation with Multimodal Large Language Model
We annotate the street view imagery using OpenAI’s ChatGPT-4o, a state-of-the-art language model designed to generate human-like text based on tailored prompts(openai2023chatgpt, ). To ensure stability and efficiency in API requests, we implement a robust, multi-threaded framework capable of managing multiple API keys to optimize the distribution of API calls. The prompts for annotation are meticulously developed based on urban building exterior indicators identified during our preliminary research. To maintain flexibility and adaptability in the annotation process, we incorporate customizable configurations, allowing modifications to suit specific research objectives.
@@
4. Experiment
4.1. Data
@@
We construct an Urban Building Exteriors Database using Street View Imagery, theoretically applicable to any urban area with street views. For this project, we select three representative cities: New York City (NYC), Amsterdam, and Singapore. The distribution of data points for the three cities is shown in Figure3.
Figure 3. The Distribution of Sampling Points
[…]
The Distribution of Sampling Points.
@@
Each city is chosen based on its unique architectural characteristics and urban landscape. New York City is renowned for its iconic skyscrapers and diverse architectural styles, ranging from historic brownstones to modern glass towers(lehan2023city, ). Amsterdam offers a distinct contrast with its picturesque canal houses, narrow buildings, and intricate facades, reflecting a historic and compact urban design(havinga2020heritage, ). Singapore, known for its modernist high-rise buildings and innovative architectural solutions, represents a dynamic and evolving urban environment that blends tradition with cutting-edge design(yan2022data, ).
@@
Figure 4. Representative Street View Images from Selected Cities
[…]
This figure showcases representative street view images of building exteriors from New York, Amsterdam, and Singapore, highlighting the architectural diversity and urban environments captured in the dataset.
@@
To collect building information, we employ the Overpass API, a powerful tool for querying OpenStreetMap(bennett2010openstreetmap, ) data, which enables us to compile comprehensive datasets for each city. We gather information on all fetched buildings within the three cities, totaling 1,382,71613827161,382,7161 , 382 , 716 buildings in NYC, 247,456247456247,456247 , 456 buildings in Amsterdam, and 298,787298787298,787298 , 787 buildings in Singapore. From these, we sample buildings and collect their building IDs, latitude and longitude coordinates, and address information. For the SVI, we utilize Google Street View Imagery(anguelov2010google, ) to capture panoramic images of individual building exteriors. To standardize the collection process, we center the building ID coordinates within a 30-meter buffer and select the closest street view image to each building. This method ensures consistency and precision in capturing building exteriors. The dataset comprises 8,13081308,1308 , 130 building images for NYC, 6,42264226,4226 , 422 building images for Singapore, and 7,75877587,7587 , 758 building images for Amsterdam. Additionally, sample street view images for each city have been included to provide visual context for the building exterior data collected, as shown in Figure 4.
4.2. Result and Analysis
@@
We gather building exterior datasets from Amsterdam (7,75877587,7587 , 758), New York City (8,13081308,1308 , 130), and Singapore (6,42264226,4226 , 422), totaling 22,3102231022,31022 , 310 records. For most metrics, the effective generation rate exceeds 99.90%. However, the effective generation rate for streetscape perception scores is comparatively lower at 86.05%, primarily due to the inherent randomness in ChatGPT’s generation process and the complexities involved in regular expression matching.
@@
Table 2. Performance Metrics for Different Variables
@@
In the accompanying figures, we select the following indicators from three perspectives—environment, energy, and human—to visualize the results: ’A’ represents the parking lot scale, ’B’ represents the tree coverage scale, and ’C’ represents the Window-to-Wall Ratio (WWR) scale, while the numbers 1, 2, and 3 correspond to Amsterdam, Singapore, and New York City, respectively. In Figure5, the color of the dots transitions from red to green, indicating the scale of the metrics from small to large. Deeper red signifies a smaller scale, while deeper green indicates a larger scale. The figure reveals that the parking lot scale has been most densely represented in red, particularly in the city centers; the tree coverage scale has been predominantly green, with darker hues; and the WWR scale has had the most intense green distribution.
@@
Figure 5. The Visualization of Result: The Number of Parking Lots (A), Trees (b) and Window-to-Wall Ratio (c)
@@
The Number of Parking Lots (A), Trees(b) and Window-to-Wall Ratio(c)
From the following word cloud, it is evident that the predominant architectural styles vary significantly across different cities. For example, in Figures A and C, ”somewhat modern” and ”modern somewhat” have been the most prevalent, whereas in Figure B, ”somewhat historical” dominates, with a relatively smaller presence in Figures A and C (see Figure6).
Figure 6. The Visualization of Result: Human Style Wordclouds
[…]
Human Style Wordclouds
5. Discussion
The building’s exterior has garnered considerable attention among researchers in the GIS domain and those engaged in urban and architectural studies. Concurrently, an increasing array of interdisciplinary professionals is examining the multifaceted influence that buildings exert on human life, the natural environment, and energy consumption. Conventional spatial databases often rely on survey data, with building attribute information typically confined to specific domains. Our conviction is that the academic community requires a comprehensive, global spatial database that documents the appearance of buildings, encompassing a wide array of significant locations.
Based on our literature review of building exteriors, we have categorized building exterior indicators across human-centric, energy, environmental, and basic information of buildings and initiated a spatial database that covers these indicators using the Large Language Model and computer vision. The characterization of building exteriors in our spatial database is grounded in the interpretations provided by ChatGPT-4o. While ChatGPT-4o’s proficiency in image comprehension is widely recognized, the absence of innate knowledge necessitates a methodical approach. To this end, we randomly select $200200$ street-level images from our database and undertake a meticulous manual review to ensure the precise depiction of building exteriors. Given that the human-centric assessments of building exteriors are inherently subjective, we have forgone verifying human-centric indicators.
We have identified and scrutinized four key indicators to evaluate the fidelity of our data: floor-to-floor height, window-to-wall ratio (WWR), the architectural vertices of the buildings, and the number of surrounding trees. Our manual review Table 2 reveals that the vision-based experiment has yielded robust results. Despite the challenges posed by occlusions inherent in street views, ChatGPT-4o has shown remarkable accuracy in correlating building heights, extrapolating door and window features, and assessing the immediate built environment. However, the number of building vertices—a three-dimensional characteristic—proved to be less accurate when relying solely on a single street view perspective. This limitation is attributed to the depth of field constraints, which can obscure the full dimensionality of the structures in question.
Despite the availability of numerous methods to construct a comprehensive global architectural spatial database, the vision-based approach significantly reduces the expenditure of human and material resources. Our experiments, while promising, reveal areas for enhancement:
  1. (1)
    The absence of street views in certain locales indicates that the creation of a complete global database of building appearances should integrate additional data sources, such as satellite imagery and expert survey data. Future endeavours could focus on developing a multi-source spatial database for building exteriors.
  2. (2)
    The single-view street view imagery presents challenges with occlusion and the limitations of a single perspective, affecting the accuracy of chatGPT-4o’s estimation of 3D building exterior attributes. Future improvements can involve leveraging images from various perspectives to generate more precise exterior descriptions.
  3. (3)
    Our current classification of building exteriors is subjective based on research trends. Moving forward, we aim to employ a more systematic text processing methodology, including optimizing chatGPT prompts, utilizing advanced NLP for text categorization, and incorporating additional databases to refine the accuracy of building exterior descriptions.
6. Conclusion
In this study, we have compiled a list of hot topics related to building exteriors for the last five years. Twenty-six metrics were summarised and categorised into disciplines. Our study presents a novel spatial database for building exteriors, demonstrating the feasibility of integrating AI and street view imagery to enhance the completeness of the database. The validation of our approach highlights its potential for efficiency and accuracy, setting a precedent for similar endeavours. Such a workflow can be replicated for global regions, resulting in a worldwide database of building exteriors.

### v2

@@ 3.3 Annotation with Multimodal Large Language Model
We annotate the street view imagery using OpenAI’s ChatGPT-4o, a state-of-the-art language model designed to generate human-like text based on tailored prompts[26]. To ensure stability and efficiency in API requests, we implement a robust, multi-threaded framework capable of managing multiple API keys to optimize the distribution of API calls. The prompts for annotation are meticulously developed based on urban building exterior indicators identified during our preliminary research. To maintain flexibility and adaptability in the annotation process, we incorporate customizable configurations, allowing modifications to suit specific research objectives.
@@
4 Experiment
This section describes the experimental setup, including data selection, sampling strategy, and validation of the proposed approach. Subsection 4.1 details the data collection from Amsterdam, New York City, and Singapore, along with the associated building and Street View datasets. Subsection 4.2 outlines the result analysis and evaluates the performance of our indicators using both manual inspection and quantitative metrics. By examining these diverse case studies, we highlight the strengths and limitations of our method.
4.1 Study Area and Dataset
@@
We construct an Urban Building Exteriors Database using Street View Imagery, theoretically applicable to any urban area with street views. For this project, we select three representative cities: New York City (NYC), Amsterdam, and Singapore. The distribution of data points for the three cities is shown in Fig3.
Figure 3: The Distribution of Sampling Points
@@
Each city is chosen based on its unique architectural characteristics and urban landscape. New York City is renowned for its iconic skyscrapers and diverse architectural styles, ranging from historic brownstones to modern glass towers[27]. Amsterdam offers a distinct contrast with its picturesque canal houses, narrow buildings, and intricate facades, reflecting a historic and compact urban design[28]. Singapore, known for its modernist high-rise buildings and innovative architectural solutions, represents a dynamic and evolving urban environment that blends tradition with cutting-edge design[29].
@@
Figure 4: Representative Street View Images from Selected Cities
@@
To collect building information, we employ the Overpass API, a powerful tool for querying OpenStreetMap[30] data, which enables us to compile comprehensive datasets for each city. From these, we sample buildings and collect their building IDs, latitude and longitude coordinates, and address information. For the SVI, we utilize Google Street View Imagery[31] to capture panoramic images of individual building exteriors. To standardize the collection process, we center the building ID coordinates within a 30-meter buffer and select the closest street view image to each building. This method ensures consistency and precision in capturing building exteriors. The dataset comprises 8,13081308,1308 , 130 building images for NYC, 6,42264226,4226 , 422 building images for Singapore, and 7,75877587,7587 , 758 building images for Amsterdam. Additionally, sample street view images for each city have been included to provide visual context for the building exterior data collected, as shown in Fig 4.
4.2 Result and Analysis
@@
We collected building exterior datasets from Amsterdam, New York City, and Singapore, comprising a total of 22,3102231022,31022 , 310 records. For most extracted indicators, the effective generation rate exceeds 99.90%. However, the generation rate for streetscape perception scores is relatively lower at 86.05%. This reduced rate is attributed to the inherent randomness in ChatGPT’s generation process and the challenges associated with complex regular expression matching.
In terms of efficiency, the entire GPT4O-based annotation process across all records incurred a total cost of approximately \$$220220$, while covering $2626$ indicators per building. This breaks down to roughly \$$0.010.01$ per building annotation. Each API call required only a few seconds on average. By deploying parallel processing, we managed to complete the entire annotation task within several hours. These results not only highlight the cost-effectiveness of our approach but also demonstrate its potential for scaling to large datasets in urban building data analysis.
@@
Table 2: Performance Metrics for Different Variables
@@
In the accompanying figures, we select the following indicators from three perspectives—environment, energy, and human—to visualize the results: ’A’ represents the parking lot scale, ’B’ represents the tree coverage scale, and ’C’ represents the Window-to-Wall Ratio (WWR) scale, while the numbers 1, 2, and 3 correspond to Amsterdam, Singapore, and New York City, respectively. In Fig5, the color of the dots transitions from red to green, indicating the scale of the metrics from small to large. Deeper red signifies a smaller scale, while deeper green indicates a larger scale.
@@
Figure 5: Visualization of Result: The Number of Parking Lots (A), Trees (b) and Window-to-Wall Ratio (c)
For parking lot scale (A1, A2, A3), deeper red areas in city centers typically indicate smaller or fewer parking lots due to higher population densities and constrained land use, while peripheral or suburban districts appear in lighter or deeper shades of green, suggesting larger or more abundant parking facilities. Tree coverage (B1, B2, B3) shows predominantly green hues, with darker greens indicating denser or more extensive tree presence. Amsterdam maintains balanced tree distribution across its canal networks and suburban edges, Singapore showcases dense greenery in residential neighborhoods (aligned with its garden city initiative), and New York City presents significant variation among boroughs, with some zones rich in street trees and others more built-up. For Window-to-Wall Ratio (C1, C2, C3), intense green signifies high ratios of windows to wall surfaces, commonly found in modern or renovated buildings favoring natural light and large windowpanes, while historic districts display red or lighter green hues due to smaller or fewer windows reflective of older architectural styles.
We also create word clouds that capture the diversity of architectural styles across the studied cities. The word clouds reveal that modern and somewhat modern architectural styles are common in some regions, whereas somewhat historical styles prevail in others (see Fig6).
The word cloud reveals notable differences in architectural styles across the three cities. In Figures A and C, the terms “somewhat modern” and “modern somewhat” are most prominent, indicating a widespread presence of contemporary architectural designs. This suggests ongoing urban development and renovation activities, which contribute to a modernized cityscape. These areas likely feature more recent constructions with streamlined facades, larger windows, and minimalist aesthetics. In contrast, Figure B shows a strong prevalence of “somewhat historical,” pointing to a significant influence of heritage architecture and preservation efforts. This pattern suggests that historical conservation policies or cultural preferences for traditional styles are more influential in this region. Consequently, the urban landscape here is characterized by older buildings with ornate facades, smaller windows, and classic architectural elements.
Additionally, the term “significant style” appears across all three maps, though its relative prominence varies. In Figures A and C, it suggests distinctive design elements that enhance visual appeal and architectural identity. In Figure B, however, its presence is more subdued, possibly indicating a more uniform historical architectural style. This variation reflects the interplay between modern development and historical preservation across the cities. The word clouds not only highlight stylistic differences but also reveal underlying cultural, historical, and socioeconomic factors that shape urban architectural identity. The diversity in architectural styles underscores how urban planning policies, cultural heritage values, and contemporary design trends contribute to the distinct character of each city.
Figure 6: The Visualization of Result: Human Style Wordclouds
5 Discussion
BuildingView provides an automated, scalable, and open-source framework for extracting building-facade attributes and contextual features from street-view imagery. In what follows, we discuss its value for various fields of built environment research in subsection 5.1, outline the role of sustainability and crowd intelligence in dataset maintenance and enhancement in subsection 5.2, and consider future directions and implementation strategies to further expand its scope and applicability in subsection 5.3.
5.1 Potential Benefits to Existing Built Environment Research
Human-Centered Urban Research
Current studies on walkability, aesthetics, and perceived safety are often limited by labor-intensive data collection, reducing research scope and comparability. BuildingView streamlines this by automatically extracting attributes like façade articulation, greenery, window transparency, and pedestrian-scale infrastructure from street-view imagery. These details provide insights into how architectural and environmental elements influence human experiences, such as neighborhood safety and public space quality, supporting evidence-based urban design to enhance walkability and visual appeal.
Scalability and Standardization in Built Environment Data
Built-environment research often lacks standardized, large-scale datasets that capture architectural complexity across different contexts. BuildingView addresses this by systematically extracting and labeling features from street-view images using fine-tuned large language models, ensuring consistent data across regions. Its open-source and modular design allows users to add new indicators or customize the framework, facilitating comparative studies and promoting standardized urban data analytics.
Urban Morphology
Traditional urban morphology studies rely on simplified datasets or field surveys with limited building details. BuildingView enhances this by using street-view imagery to capture detailed façade geometry, window-to-wall ratios, and architectural elements, offering high-resolution morphological indicators. These observations reveal architectural patterns, urban density impacts on livability, and adaptations to cultural or climatic contexts, enabling precise cross-city morphological comparisons and better classification of architectural styles.
5.2 Sustainability and Crowd Intelligence
BuildingView is fully automated and reproducible, ensuring continuous updates and responsiveness to urban changes. Scripts for data acquisition, annotation, and enrichment are openly available, enabling scheduled re-runs when new street-view imagery becomes available. As urban environments evolve—new buildings emerge, façades are updated, or land-use patterns change—these automated updates keep the data relevant. Future improvements can integrate advanced building attribute prediction models for greater accuracy and richer metadata.
In addition to automated updates, BuildingView leverages crowd intelligence, turning data users into contributors. In areas with limited street-view coverage, local communities or researchers can enhance the dataset using open mapping platforms like Mapillary or KartaView. An interactive web-based validation tool could also allow volunteers to verify or refine architectural labels, ensuring higher data quality through community participation. This collaborative, open-source approach supports global open-data principles, making BuildingView a valuable resource for urban researchers, planners, and policymakers.
5.3 Future Directions and Implementation Considerations
BuildingView is designed to accommodate a variety of urban contexts and building types using street-view data. Ongoing and future enhancements can amplify its robustness, usability, and overall impact. First, broadening geographic reach remains a key priority for capturing buildings in regions where street-view coverage is limited or inconsistent, a process that may benefit from partnerships with local institutions or community-driven imagery initiatives. Second, refining classification algorithms and annotation prompts can help distinguish subtle features and address any class imbalances, especially for rare building elements. Third, volunteer-driven validation tools could allow local users to verify automated labels in challenging scenarios, strengthening data reliability and inclusiveness. Lastly, reducing reliance on proprietary APIs can facilitate more equitable access and sustainability by leveraging open-source or collaborative data platforms where feasible. Through these incremental improvements, BuildingView can serve as a valuable resource for built environment research, policy-making, and community engagement, supporting a clearer understanding of how façades and streetscapes shape the social, environmental, and energy dimensions of cities.
6 Conclusion
In this study, we have compiled a list of hot topics related to building exteriors for the last five years. Twenty-six metrics are summarised and categorized into disciplines. Our study presents a novel spatial database for building exteriors, demonstrating the feasibility of integrating AI and street view imagery to enhance the completeness of the database. The validation of our approach highlights its potential for efficiency and accuracy, setting a precedent for similar endeavors. Such a workflow can be replicated for global regions, resulting in a worldwide database of building exteriors.
{credits}
6.0.1 Acknowledgements
This work is supported by the National Natural Science Foundation of China (Grant No. 42401567), the Tertiary Education Scientific Research Project of the Guangzhou Municipal Education Bureau (Grant No. 2024312159), and the Guangzhou Municipal Science and Technology Bureau Program (Grant No. 2025A03J3640).
@@ 6.0.1 Acknowledgements
The authors have no competing interests to declare that are relevant to the content of this article.